### v1

# Lavida-O: Elastic Masked Diffusion Models for Unified Multimodal Understanding and Generation

Abstract
We proposed Lavida-O, a unified multi-modal Masked Diffusion Model (MDM) capable of image understanding and generation tasks. Unlike existing multimodal diffsion language models such as MMaDa and Muddit which only support simple image-level understanding tasks and low-resolution image generation, Lavida-O exhibits many new capabilities such as object grounding, image-editing, and high-resolution (1024px) image synthesis. It is also the first unified MDM that uses its understanding capabilities to improve image generation and editing results through planning and iterative self-reflection. To allow effective and efficient training and sampling, Lavida-O introduces many novel techniques such as Elastic Mixture-of-Transformer architecture, universal text conditioning, and stratified sampling. Lavida-O achieves state-of-the-art performance on a wide range of benchmarks such as RefCOCO object grounding, GenEval text-to-image generation, and ImgEdit image editing, outperforming existing autoregressive and continuous diffusion models such as Qwen2.5-VL and FluxKontext-dev, while offering considerable speedup at inference.
Figure 1: We propose Lavida-O, a unified masked diffusion model capable of multi-modal understanding and generation.
1 Introduction
The abilities to understand and generate images have been two essential objectives of image modeling research. Traditionally, these tasks are handled by a diverse set of specialist models, such as detection models for object localization (Liu et al., 2024b; Li et al., 2023a), Visual Question Answering (VQA) models for question-answering (Li et al., 2022), and diffusion models for text-to-image generation (Esser et al., 2024; Podell et al., 2023; Rombach et al., 2022). Recently, the rise of unified multi-modal models such as GPT-4o (OpenAI, 2024) has introduced a new paradigm: using a single generalist model to perform a wide range of image understanding and generation tasks. Not only is this unified approach more aligned with the goal of developing versatile multi-task Artificial General Intelligence (AGI), but it also demonstrates strong empirical performance by allowing understanding and generation capabilities to mutually benefit each other (Deng et al., 2025). This is especially notable in tasks requiring both understanding and generation capabilities, such as image editing, where unified models show unparalleled advantages over generation specialists.
Most current unified models are built on Autoregressive (AR) large language models. Some works, such as BLIP3o (Chen et al., 2025a) and BAGEL (Deng et al., 2025), employ AR modeling for text generation and continuous diffusion modeling for image generation (AR+diff), while others, such as Janus (Chen et al., 2025c), first tokenize images into sequences of discrete tokens and then employ a unified AR next-token prediction objective for both image and text modalities.
Recently, Masked Diffusion Models (MDMs) (Lou et al., 2023; Sahoo et al., 2024) have emerged as a competitive alternative to AR models. Unlike AR models, MDMs treat token generation as a diffusion process over discrete tokens. In the forward process, the tokens of a sequence are gradually masked. At inference, we start with a sequence of mask tokens and gradually unmask them to obtain a sequence of meaningful tokens. Large-scale experiments in language modeling (Nie et al., 2025; Ye et al., 2025a) show that MDMs can achieve comparable performance to AR language models while offering many advantages, such as better speed-quality tradeoffs, controllability, and bi-directional context. Several recent works extend MDMs to multi-modal understanding and generation tasks (Li et al., 2025a; Yu et al., 2025b; Yang et al., 2025; Shi et al., 2025). Compared with the AR+diff setup, unified MDMs avoid the need to carefully tune the balance between AR and diffusion losses by offering a unified objective, resulting in greater simplicity and scalability. Compared with unified AR modeling, unified MDMs offer significantly faster sampling speeds by allowing parallel decoding of multiple tokens.
Despite these advantages, the latest unified MDMs—such as MMaDa (Yang et al., 2025) and Muddit (Shi et al., 2025)—still lag behind state-of-the-art unified AR and AR+diffusion models, both in the breadth of tasks they support and in benchmark performance. There are three main challenges in developing high-performing unified MDMs. First, unified models are expensive to train due to the large size of their language backbones. For example, to build a unified MDM with image generation capability, MMaDa pretrains an 8B model jointly on text and image generation, which is costly. This challenge is further exacerbated by the limited literature on training large-scale masked image generative models. In contrast, many open-sourced large-scale continuous diffusion models such as Flux (Labs, 2024) are readily available. Second, open-sourced resources for masked image generative models (MIGMs) are limited, and the literature on their training techniques and sampling processes is less developed than that for continuous diffusion models. Even the best open-sourced MIGM, Meissonic-1B (Bai et al., 2024), significantly underperforms continuous diffusion models of comparable size (Xie et al., 2025a). Lastly, while these models can perform both understanding and generation tasks, they lack explicit mechanisms to leverage image understanding capabilities to improve generation quality. In fact, MMaDa and Muddit cannot even perform image editing tasks, which require both understanding and generation capabilities. These models simply concatenate text-to-image data and image understanding data during training.
To bridge this gap, we propose Lavida-O, a unified multi-modal Masked Diffusion Model (MDM) capable of both image understanding and generation tasks. To mitigate the cost of training large diffusion models, Lavida-O introduces several techniques such as Elastic Mixture-of-Transformers (Elastic-MoT), progressive upscaling (gradually increasing the image resolution during training), and token compression that enable efficient scaling. To improve generation quality, Lavida-O employs stratified sampling and universal text conditioning. To fully leverage the potential of a unified multi-modal model, Lavida-O incorporates planning and self-reflection mechanisms that explicitly utilize its understanding capabilities to enhance generation outputs. We highlight Lavida-O’s capabilities compared with previous multi-modal MDMs in Table 1.
[…]
  • We propose the first multi-modal MDM that achieves state-of-the-art performance on text-to-image generation, image editing, and grounding tasks, outperforming existing MDMs, AR models, and continuous diffusion models.
  • We propose several efficient and effective training and inference techniques for large-scale masked image generative models and unified multi-modal models, such as the Elastic-MoT architecture, universal text conditioning, and stratified sampling, significantly advancing the literature.
  • We introduce a novel paradigm that explicitly leverages the understanding capabilities of a unified model to improve its generation through planning and self-reflection.
Table 1: Capabilities of different multimodal MDMs. Lavida-O uniquely supports localized understanding, high-resolution image synthesis, image editing and interleaved generation. *Muddit showed examples of simple editing through inpainting. It does not have instruction-based editing capabilities.
Understanding Generation
Model Image-level Object-level Text-to-image Image-editing Interleaved
LaViDa $\checkmark$ $\times$ $\times$ $\times$ $\times$
Dimple $\checkmark$ $\times$ $\times$ $\times$ $\times$
Muddit $\checkmark$ $\times$ $512^{2}$ * $\times$
MMaDa $\checkmark$ $\times$ $512^{2}$ $\times$ $\times$
LaViDa-O $\checkmark$ $\checkmark$ $1024^{2}$ $\checkmark$ $\checkmark$
2 Background and Related Works
2.1 Masked Diffusion Models
Masked Generative Modeling (MGM) has emerged as an alternative to AR models for modeling sequences of discrete tokens. In MGM, sequences are partially masked during training, and the model learns to predict the masked tokens. Early works such as BERT (Devlin et al., 2019) used MGM as a representation learning objective to extract semantically rich features for downstream tasks. Later works like MaskGIT (Chang et al., 2022) and MUSE (Chang et al., 2023) explored using MGM for generative modeling. In this setup, a sequence is initialized with only mask tokens, which are gradually unmasked by a neural network during sampling to generate the desired output. These works primarily focus on image generation tasks, where discrete image tokenizers such as VQGAN (Esser et al., 2021) are used to convert images into sequences of discrete tokens.
More recently,MDMs (Austin et al., 2021; Sahoo et al., 2024; Lou et al., 2023) have further developed the theory of MGM by formalizing the masking and unmasking process as the forward and reverse diffusion processes in discrete space. This provides a principled framework for training and sampling from these models. MDMs have renewed interest in masked modeling for language generation, offering theoretical advantages over AR models, such as better speed-quality tradeoffs and improved controllability. Notably, LLaDa-8B and Dream-8B (Nie et al., 2025; Ye et al., 2025a) demonstrated that MDMs can achieve competitive performance compared to AR models at scale. Several follow-up works (Li et al., 2025a; Yu et al., 2025b; You et al., 2025; Yang et al., 2025), including LaViDa and MMaDa, extend MDMs to multi-modal tasks such as image understanding and text-to-image generation. Their capabilities are summarized in Table 1.
Formally, given a sequence of $L$ discrete tokens $X_{0}=[X_{0}^{1},X_{0}^{2},\dots,X_{0}^{L}]$, the forward process $q(X_{t}|X_{s})$ gradually masks the tokens over the time interval [0,1], with $1\geq t\geq s\geq 0$. At $t=1$, the sequence $X_{1}=[X_{1}^{1},X_{1}^{2},\dots,X_{1}^{L}]$ consists entirely of masked tokens, denoted by $[M]$. A neural network $p_{\theta}$ is used to model the reverse process $p(X_{s}|X_{t})$. The masked diffusion objective is defined as:
$\mathcal{L}_{\text{MDM}}=-\mathbb{E}_{t,X_{0},X_{t}}\left[\frac{1}{t}\log p_{\theta}(X_{0}|X_{t})\right]=-\mathbb{E}_{t,X_{0},X_{t}}\left[\frac{1}{t}\sum_{i\in\{1,2,\dots,L\},X_{t}^{i}=[M]}\log p_{\theta}(X_{0}^{i}|X_{t})\right]$ (1)
where $\log p_{\theta}(X_{0}|X_{t})$ is factorized into $\prod_{i=1}^{L}p_{\theta}(X_{0}^{i}|X_{t})$ based on independence assumptions (Sahoo et al., 2024). At inference time, the model starts from a fully masked sequence $X_{1}=[M,M,\dots,M]$ and progressively applies the learned reverse process $\log p_{\theta}(X_{0}|X_{t})$ to recover the original tokens. We provide a more detailed formulation of MDMs in Appendix A.1.
2.2 Unified Multi-modal Models
Unified multi-modal models such as GPT-4o (OpenAI, 2024) are capable of both image understanding and generation tasks, leading to strong performance on tasks requiring both capabilities, such as image editing. Generally, there are two dominant types of unified models based on their modeling objectives. The first type employs an autoregressive next-token prediction objective for text generation and a diffusion objective for image generation (AR+diff). Notable models in this category include BLIP-3o (Chen et al., 2025a), BAGEL (Deng et al., 2025), and Metaqueries (Pan et al., 2025). However, this design involves two different training objectives with distinct numerical scales and training dynamics, often requiring careful tuning of loss weighting and data mixtures.
In contrast, the second type of models employ a unified objectives for both image and texts. Early works like Janus-Pro (Chen et al., 2025c) employ a unified AR modeling objective. Several recent works like MMaDa (Yang et al., 2025) explore a unified masked diffusion objective. Despite some success, a significant performance gap remains between these unified MDMs and state-of-the-art unified models in the AR and AR+diff categories.
Architecturally, unified models also fall into two main categories. The first type, such as Janus and MMaDa, uses a single dense transformer to output both image and text tokens. The second type, such as BAGEL and MetaQueries, employs separate parameter sets for handling image and text modalities. A common design in this category is the mixture-of-transformers (MoT) architecture (Liang et al., 2024), where image and text inputs are processed by different parameter sets but can interact through joint attention mechanisms. These designs are illustrated in Figure 3. While being more flexible, training MoT experts can be expensive due to the duplication of parameters. For example, BAGEL extends a 7B Qwen model into a 7B+7B MoT, doubling the parameter count.
3 Method
Figure 2: Overall Pipeline of Lavida-O. Given an input image and text prompt, we first concatenate the image semantic embedding $C_{i}$, image VQ embedding $C_{v}$, and text prompt embedding $C_{p}$ to form the conditioning embedding $C$. The combined embedding is then passed to the model alongside the partially masked sequence $X_{t}$. The model then predicts the fully-unmasked sequence $X_{0}$.
3.1 Model Architecture
Lavida-O’s model architecture is built on LaViDa (Li et al., 2025a), a diffusion vision-language model capable of image understanding tasks. LaViDa uses a SigLIP (Zhai et al., 2023) vision encoder to convert input images into continuous semantic embeddings $C_{i}$, which are concatenated with token embeddings of text prompts $C_{t}$ to form the final conditional embeddings $C=\text{Concat}(C_{i},C_{p})$ for visual understanding tasks. At each inference step, the diffusion model uses the partially unmasked answer $X_{t}$ and the conditional embedding $C$ to predict the reverse distribution $p_{\theta}(X_{0}|X_{t},C)$, where $X_{0}$ is the clean text answer.
For image understanding tasks, Lavida-O maintains this eact setup of LaViDa. To incorporate visual generation tasks, we extend LaViDa’s design by representing target images as sequences of discrete tokens using a VQ-Encoder (Esser et al., 2021). When performing these tasks, $X_{0}$ and $X_{t}$ contain not only text tokens, but also VQ tokens representing images. For image editing and interleaved generation tasks, we additionally incorporate VQ tokens of input images $C_{v}$ as part of the conditional embedding $C=\text{Concat}(C_{i},C_{v},C_{p})$, since using semantic embeddings $C_{i}$ alone may lose low-level details needed for editing. To reduce the number of tokens and improve computational efficiency, we introduce a token compression module that reduces the number of VQ tokens by a factor of 4. The overall pipeline is illustrated in Figure 2.
3.1.1 Elastic Mixture-of-Transformers (ElasticMoT)
Our goal is to find an effective and efficient method that can equip LaViDa, an understanding-only diffusion model, with visual generation capabilities. However, both of the existing common choices described in Section 2.2—dense models and MoT—are very expensive. Dense models use the same set of parameters for both understanding and generation tasks, requiring a mix of understanding and generation data during training to prevent catastrophic forgetting, which is not data-efficient. While the MoT setup allows freezing the understanding branch and training only the generation branch for image generation, its architecture doubles the total parameter count, leading to considerable computational overhead. Moreover, given an 8B base understanding model, both setups require training at least 8B parameters for generation tasks from scratch, which is prohibitively expensive.
To address these limitations, we propose Elastic-MoT, a novel architecture design that efficiently adapts an understanding-only model for image generation tasks. Compared with the vanilla MoT architecture, Elastic-MoT introduces two major modifications. First, instead of using equally sized understanding and generation branches, the generation branch has a smaller hidden size. This reduces the parameter count and enables efficient training. We make this design choice based on the observation that many text-to-image models can generate high-quality images with only 2–4B parameters, suggesting that generation tasks may not require as much capacity as understanding tasks (Xie et al., 2025a, b).
Second, given an $N$-layer model, instead of having joint attention at all layers, we only allow text and image modalities to interact in the first $M$ layers. In the remaining K=NMK=N-M layers, text and image tokens interact only within their modality through self-attention. This design activates only partial parameters for different tasks. For example, in Lavida-O’s final design, the generation branch consists of 2.4B new parameters, while the understanding branch has 8B parameters initialized from LaViDa. With $N=32$ layers, we set $M=K=16$. For image generation tasks, we activate all 2.4B parameters in the image generation branch and 4B parameters from the first 16 layers of the understanding branch, leading to 6.4B active parameters. During text-to-image pretraining, we train only the 2.4B parameters in the generation branch, and omit loading the last 16 layers of the understanding branch since they are unused. This enables efficient scaling of text-to-image pretraining without introducing extra understanding data or bearing the large parameter count of MoT. Similarly, the active parameter count is 8B for understanding tasks, and 10.4B for tasks requiring both understanding and generation, such as interleaved generation. The full design of Elastic-MoT is illustrated in Figure 3. We provide additional details and analysis in Appendix A.2 and B.2.
Figure 3: Design of Elastic MoT. Elastic-MoT introduces two major modifications to standard MoT. First, the generation branch has a smaller hidden size. Second, given an $N$-layer model, we only allow text and image modalities to interact in the first $M$ layers. These two designs allow us to flexibly load only a portion of parameters depending on tasks, improving the efficiency.
3.1.2 Modality-aware Masking
One of the challenges in adapting MoT architecture for MDMs is routing—the mechanism to determine which branch should be activated for each token. This is trivial for unified AR MoT models, where the model can simply learn to generate a special token (e.g., [img_start]) to indicate that the next token should use the generation branch. When the last token of an image is generated, the model automatically switches back to the understanding branch to continue text generation. However, this approach is not feasible for MDMs, which decode tokens in parallel and must decide in advance which mask tokens should be routed to the understanding branch and which to the generation branch. A naive solution is to let the user specify the number and location of text and image tokens, but this is difficult for interleaved generation, such as image generation with self-reflection. To address this issue, we design a modality-aware masking process.
Given a sequence of $M$ text tokens and $N$ image VQ tokens, the vanilla forward diffusion process gradually converts it into $M+N$ mask tokens during the time interval $[0,1]$. By contrast, our modality-aware forward process introduces a special timestamp $t_{\text{exp}}\in[0,1]$, at which fully masked image VQ tokens are collapsed into a special $\text{exp}]$ text token. This process is illustrated in Figure 4(a) (Bottom-up). At inference, we assume all mask tokens are text tokens at the beginning. When a $\text{exp}]$ token is generated, we replace it with a sequence of $L_{\text{img}}$ mask tokens, and specify that these tokens will be processed by the generation branch for image synthesis in subsequent forward calls. $L_{\text{img}}$ is determined by a pre-specified image resolution. This process is also illustrated in Figure 4(a) (Top-down). We provide additional details in Appendix A.3.
(a) Modality-Aware Masking
(b) Stratified Sampling
Figure 4: Design choices of Lavida-O. (a) Forward diffusion process with modality-aware masking. (b) Visualization of the unmasking order in the proposed stratified random sampling process.
3.2 Task-Specific Designs
In this section, we describe several additional technical innovations that improve the effectiveness and efficiency on newly incorporated tasks such as image generation, image editing and grounding.
Universal Text Conditioning. A common approach to improving the quality of text-to-image models is micro-conditioning (Bai et al., 2024; Podell et al., 2023), which conditions the image generation process on extra parameters such as original image resolution, crop coordinates, and image quality scores. This is typically achieved via specialized embeddings or modulation layers. However, since a unified model has strong language understanding and reasoning capabilities, we can simply append these conditions as plain text (e.g., “SCORE: 5.40”) to the end of user prompts. In addition to common conditions, we also incorporate image luminance and contrast as micro-conditions. This simple and effective design not only improves image quality by biasing generation toward high-scoring distributions, but also gives users more refined control over outputs. We provide additional details in Appendix A.4.
Stratified Random Sampling. Existing MDMs mostly adopt confidence-based sampling at inference, where tokens with high confidence scores are unmasked first. However, this design is suboptimal for image generation, where many tokens need to be decoded at each step. In particular, high-confidence tokens tend to cluster in small spatial regions adjacent to already unmasked tokens. This negatively affects image quality, since adjacent tokens are highly correlated, yet MDMs sample all unmasked tokens independently and assume their joint distribution perfectly factorizes. To mitigate this, we aim to minimize correlation between unmasked tokens at each step by maximizing their spatial distance. We achieve this with a stratified random sampling process, which ensures unmasked tokens are evenly distributed across the image. Specifically, we first divide the image into $2\times 2$ square regions and unmask one random token per region. We then subdivide each region into 4 subregions, giving 16 in total. Four already contain an unmasked token; we randomly unmask one token from each of the remaining 12. We repeat this process until all tokens are unmasked. This is illustrated in Figure 4(b). More details and analysis are provided in Appendix A.5 and B.3.
Planning and Reasoning. While existing unified MDMs integrate image understanding and generation tasks with a single diffusion objective, they do not incorporate mechanisms that use understanding to improve generation, except for the assumption that joint training benefits both tasks. To address this, we introduce two explicit mechanisms that leverage understanding to improve generation: planning and reflection. With planning, the model first generates a layout of the image represented by bounding boxes, then creates the actual image accordingly. For image editing tasks, it first identifies the desired edit region before generating the edited image. With reflection, the model evaluates its own generation using its understanding capability and determines whether it satisfies the user’s request. If misalignment is detected, the model generates a new image correcting the error. Examples are shown in Figure 1, with additional results in the appendix.
Object Grounding with Coordinate Quantization. The bi-directional context of MDMs naturally allows parallel decoding of bounding box coordinates. While Lavida-O can represent numbers as plain text, we adopt a specialized scheme that normalizes all bounding box coordinates to $[0,1]$ and quantizes them into 1025 discrete tokens representing $\frac{0}{1024},\frac{1}{1024},...,\frac{1024}{1024}$. This ensures each bounding box is represented by exactly four tokens. At inference, we construct a multiple query input with masked tokens such as “A dog [m][m][m][m]; A cat [m][m][m][m]”, and unmask all coordinates in parallel. This design allow us to decode multiple bounding boxes in a single diffusion step, greatly boosting the efficiency. We provide further details in Appendix A.6
4 Experiment
4.1 Setup
We start with LaViDa, an 8B understanding-only diffusion model, and extend it with a 2.4B image generation branch using the ElasticMoT architecture described in Section 3.1.1. The training consists of three stages:
  • Stage 1: We continue training the base model on object grounding and image-level understanding tasks using 12M samples. In this stage, the generation branch is not activated.
  • Stage 2: We pretrain the 2.4B image generation branch for text-to-image generation tasks using 200M image-text pairs. In this stage, we load 2.4B + 4B parameters using the ElasticMoT design described in Section 3.1.1. We freeze the understanding branch and train only the generation branch, thereby avoiding the need to include understanding data to prevent forgetting. To further improve training efficiency, we start at a resolution of 256 and progressively increase it to 512 and 1024 during training.
  • Stage 3: In the final stage, we jointly train the entire 2.4B + 8B model end-to-end on image-level understanding, text-to-image generation, image editing, and interleaved generation tasks such as planning and self-reflection. In addition to the data used in previous stages, we incorporate 1.5M image editing samples and 3M interleaved generation samples.
More details about the training data and process are provided in Apendix B.1.
4.2 Image Understanding
We report performance on image understanding tasks in Table 2. Lavida-O outperforms the previous state-of-the-art unified diffusion model, MMaDa, by a considerable margin on MMMU (Yue et al., 2024), MME (Fu et al., 2023), and MMB (Liu et al., 2024c). Compared with the base model LaViDa, Lavida-O achieves substantial improvements on most benchmarks such as ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021), ScienceQA (Lu et al., 2022), and MathVista (Lu et al., 2023), thanks to the scaling of training data.
Table 2: Quantitative results on image-level understanding tasks. *We evaluated results not reported in the original paper using official weights.
Model MMMU MME-P MME-C MMB ChartQA DocVQA InfoVQA Sci.QA AI2D M.Vista M.Verse
AR Und. Only
LLaVa-1.6-7B (Liu et al., 2024a) 35.1 1519.3 323 54.6 64.9 74.4 37.1 73.2 66.6 34.4 14.3
Qwen2.5-VL-7B (Bai et al., 2025) 58.6 - - 83.5 84.9 82.6 - - 83.9 68.2 49.2
Intern-VL-3-8B (Li et al., 2024a) 65.6 - - 83.4 86.6 92.7 76.8 - 85.2 75.2 39.8
AR Unified Und. and Gen.
BAGEL (Deng et al., 2025) 55.3 1687 701 85 - - - - - 73.1 -
Show-O (Xie et al., 2024) 27.4 1233 - - - - - - - -
Masked Und. Only
Dimple (Yu et al., 2025b) 45.2 1514 432 74.6 63.4 - - 77.1 74.4 42.3 -
LaViDa (Li et al., 2025a) 43.6 1366 341 70.5 64.6 59.0 34.2 80.2 70.0 44.8 27.2
Masked Unified Und. and Gen.
Muddit (Shi et al., 2025) - 1104 - - - - - - - - -
MMaDa (Yang et al., 2025) 30.2 1410 242* 68.5 9.8* 10.9* 14.9* 55.8* 66.6* 33.7* 13.5*
LaViDa-O 45.1 1431 488 76.4 80.0 73.7 44.6 84.6 76.7 56.9 36.9
4.3 Text-to-Image Generation
We report text-to-image generation results on the GenEval (Ghosh et al., 2023) and DPG (Hu et al., 2024) benchmarks, and FID scores on 30k prompts from MJHQ (Li et al., 2024b) dataset. We compare against text-to-image models including Flux-dev (Labs, 2024), SD3-Medium (Esser et al., 2024), Meissonic (Bai et al., 2024) and DALLE-3 (OpenAI, 2023), unified models including BAGEL (Deng et al., 2025), Show-o (Xie et al., 2024), MMaDa (Yang et al., 2025) and Muddit (Shi et al., 2025). Lavida-O significantly outperforms the state-of-the-art masked image generation model Meissonic, as well as unified models like MMaDa and Muddit. Planning and reflection further enhance prompt-following performance. We did not activate planning and reflection on MJHQ due to its large size and that FID does not reflect prompt-following capabilities.
Table 3: Quantative results on text-to-image generation tasks. *We evaluated results not reported in the original paper using official weights.
[…]
Gen. Only
Flux-dev (Labs, 2024) 12B Continuous 0.68 84.0 10.15
SD3-Medium (Esser et al., 2024) 2B Continuous 0.74 84.1 11.92
DALLE-3 (OpenAI, 2023) - Continuous 0.67 83.5 -
Meissonic (Bai et al., 2024) 1B Masked 0.54 - -
Unified Und. and Gen.
BAGEL (Deng et al., 2025) 7B+7B Continuous 0.82 - -
OmniFlow (Li et al., 2024c) 3.4B Continuous 0.62 - -
Show-o (Xie et al., 2024) 1.3B Masked 0.67 - 15.18
Muddit (Shi et al., 2025) 1B Masked 0.61 - -
MMaDA (Yang et al., 2025) 8B Masked 0.63 53.4* 32.85*
LaViDa-O 4B+2.4B Masked 0.77 81.8 6.68
+Planning 8B+2.4B Masked 0.85 82.9 -
+Reflection 8B+2.4B Masked 0.89 83.2 -
4.4 Object Grounding
We evaluate the object grounding capabilities of Lavida-O on RefCOCO Referring Expression Comprehension (REC) tasks, reporting the Precision@0.5 metric. Lavida-O outperforms autoregressive vision-language models such as Qwen2.5-VL-7B (Bai et al., 2025) and InternVL3-8B (Zhu et al., 2025), as well as specialist models such as Grounding-DINO-L (Liu et al., 2024b) and SegLLM-7B (Wang et al., 2025a).
Table 4: Quantative results on RefCOCO, RefCOCO+, and RefCOCOg REC tasks. We report Precision@0.5 on object grounding tasks and compare against baseline models.
Model RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
SegLLM-7B(Wang et al., 2025a) 90.0 92.1 86.2 82.2 85.5 76.1 83.9 85.9
Qwen2.5-VL-7B (Bai et al., 2025) 90.0 92.5 85.4 84.2 89.1 76.9 87.2 87.2
GroundingDINO (Liu et al., 2024b) 90.6 93.2 88.2 88.2 89.0 75.9 86.1 87.0
InternVL3-8B (Zhu et al., 2025) 92.5 94.6 88.0 88.2 92.5 81.8 89.6 90.0
LaViDa-O 92.3 94.8 89.0 88.7 92.5 83.3 90.0 90.6
4.5 Image Editing
We report image editing performance on the Image-Edit Bench in Table 5. Lavida-O outperforms state-of-the-art unified models such as BAGEL and specialized models like FluxKontext-dev. Most notably, Lavida-O even outperforms the state-of-the-art closed-sourced model GPT4-o(OpenAI, 2024) on replacing and removing objects, which requires localized understanding of the image. This underscores the effectiveness of Lavida-O’s design in integrating object grounding capabilities.
Table 5: Quantative results on Image-Edit benchmark. We report the per-category and overall score on Image-Edit Benchmark.
Model Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
GPT-4o (OpenAI, 2024) 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.89 4.20
Qwen2.5VL+Flux (Wang et al., 2025b) 4.07 3.79 2.04 4.13 3.89 3.90 4.84 3.04 4.52 3.80
FluxKontext dev (Labs et al., 2025) 3.76 3.45 2.15 3.98 2.94 3.78 4.38 2.96 4.26 3.52
OmniGen2 (Wu et al., 2025b) 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 4.68 3.44
UniWorld-V1 (Lin et al., 2025) 3.82 3.64 2.27 3.47 3.24 2.99 4.21 2.96 2.74 3.26
BAGEL (Deng et al., 2025) 3.56 3.31 1.70 3.30 2.62 3.24 4.49 2.38 4.17 3.20
Step1X-Edit (Liu et al., 2025) 3.88 3.14 1.76 3.40 2.41 3.16 4.63 2.64 2.52 3.06
ICEdit (Zhang et al., 2025) 3.58 3.39 1.73 3.15 2.93 3.08 3.84 2.04 3.68 3.05
OmniGen (Xiao et al., 2025) 3.47 3.04 1.71 2.94 2.43 3.21 4.19 2.24 3.38 2.96
UltraEdit (Zhao et al., 2024) 3.44 2.81 2.13 2.96 1.45 2.83 3.76 1.91 2.98 2.70
AnyEdit (Yu et al., 2025a) 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.65 2.45
InstructAny2Pix(Li et al., 2023b) 2.55 1.83 2.10 2.54 1.17 2.01 3.51 1.42 1.98 2.12
MagicBrush (Zhang et al., 2023) 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.22 1.90
Instruct-Pix2Pix(Brooks et al., 2023) 2.45 1.83 1.44 2.01 1.50 1.44 3.55 1.20 1.46 1.88
LaViDa-O 4.04 3.62 2.01 4.39 3.98 4.06 4.82 2.94 3.54 3.71
+ Planning 4.11 3.67 2.04 4.40 4.05 4.00 4.75 3.10 4.04 3.80
4.6 Training and Inference Speed
In Figure 5, we benchmark the inference efficiency of Lavida-O across three tasks: text-to-image generation, object grounding, and math reasoning. We measure end-to-end latency in seconds per image. Lavida-O is significantly faster than autoregressive models. Notably, we achieve a $6.8\times$ speedup on object grounding tasks compared to Qwen2.5-VL-7B (Bai et al., 2025). We also report the training efficiency measured by per-step latency and compare our Elastic-MoT design with BAGEL-style standard MoT design, Elastic-MoT improves the training speed by $3.17\times$. We provide additional analysis on the speed-quality tradeoff at inference time in Appendix B.6 and analysis on the training efficiency of Elastic-MoT design in B.2.
Figure 5: Training and Inference Speed of Lavida-O. We compare the end-to-end inference latency of Lavida-O on three distinct tasks: text-to-image generation, object grounding, and math reasoning. We also report the pretraining efficiency measured by per-step latency.
5 Conclusion
In summary, we proposed Lavida-O, the first multi-modal masked diffusion model that achieves state-of-the-art performance on text-to-image generation, image editing, and grounding tasks—competitive with the best specialist models and autoregressive unified models.
We also introduced a novel paradigm of interleaved generation, which explicitly leverages understanding capabilities to improve generation results in a unified multi-modal model through planning and self-reflection.
In developing Lavida-O, we proposed several efficient training and inference techniques, including the ElasticMoT architecture, universal text conditioning, and stratified random sampling, providing valuable insights for future work in masked diffusion models and unified multi-modal systems.
Appendix A Additional Technical Details
A.1 Formulation of Masked Diffusion Models
Masked Diffusion Models (MDMs) model the generation process of discrete token sequences through a continuous-time Markov chain (CMTC). Formally, given a sequence of discrete tokens $X_{0}=[X_{0}^{1},X_{0}^{2},\ldots,X_{0}^{L}]$ of length $L$, the forward process $q(X_{t}|X_{s})$ gradually converts it into a sequence of mask tokens $[M]$, denoted by $X_{1}=[X_{1}^{1},X_{1}^{2},\ldots,X_{1}^{L}]$, over the continuous time interval $[0,1]$, with $1\geq t\geq s\geq 0$. Each token $X_{t}^{i}$ belongs to a fixed-size vocabulary set $V$. In our setup, $V$ consists of text tokens, image VQ tokens, and the special mask token $[M]$. This forward process is formally defined as
$q(X_{t}^{i}|X_{s}^{i})=\begin{cases}\text{Cat}(X_{t}^{i};\textbf{M}),&\text{if }X_{s}^{i}=[M]\\ \text{Cat}(X_{t}^{i};\frac{1-t}{1-s}\mathbf{X_{s}^{i}}+\frac{t-s}{1-s}\textbf{M}),&\text{if }X_{s}^{i}\neq[M],\end{cases}$ (2)
where $\text{Cat}(\cdot)$ denotes a categorical distribution, and $\textbf{M},\mathbf{X_{0}^{i}},\mathbf{X_{s}^{i}}\in\mathbb{R}^{|V|}$ are probability vectors, with $|V|$ denoting the vocabulary size. In particular, M is a one-hot vector representing the mask token $[M]$. This forward process yields the following marginal distribution:
$q(X_{t}^{i}|X_{0}^{i})=\text{Cat}(X_{t}^{i};(1-t)\mathbf{X_{0}^{i}}+t\textbf{M}).$ (3)
MDLM (Sahoo et al., 2024) demonstrated that the posterior of the reverse process $p(X_{s}|X_{t},X_{0})$ has the following form:
$p(X_{s}^{i}|X_{t}^{i},X_{0}^{i})=\begin{cases}\text{Cat}(X_{s}^{i};\mathbf{X_{t}^{i}}),&\text{if }X_{s}^{i}\neq[M]\\ \text{Cat}(X_{s}^{i};\tfrac{t-s}{t}\mathbf{X_{0}^{i}}+\tfrac{s}{t}\textbf{M}),&\text{if }X_{s}^{i}=[M].\end{cases}$ (4)
In practice, we replace $\mathbf{X_{0}^{i}}$ with the neural network prediction $p_{\theta}(X_{0}^{i}|X_{t})$ when sampling from the reverse process, which gives the following transition:
$p_{\theta}(X_{s}^{i}|X_{t})=\begin{cases}\text{Cat}(X_{s}^{i};\mathbf{X_{t}^{i}}),&\text{if }X_{s}^{i}\neq[M]\\ \text{Cat}(X_{s}^{i};\tfrac{t-s}{t}p_{\theta}(X_{0}^{i}|X_{t})+\tfrac{s}{t}\textbf{M}),&\text{if }X_{s}^{i}=[M].\end{cases}$ (5)
Sampling process. At inference time, we initialize $X_{1}$ as a sequence of mask tokens, with $X_{1}^{1}=X_{1}^{2}=\cdots=X_{1}^{L}=[M]$. We discretize the continuous time interval $[0,1]$ into discrete timesteps $0=t_{0}<t_{1}<\cdots<t_{K}=1$, and iteratively sample $X_{t_{k-1}}\sim p_{\theta}(X_{t_{k-1}}|X_{t_{k}})$ using Equation 5. We start with $k=K$ and end when we obtain a mask-free sequence $X_{0}$. At each step, we sample each token position independently, assuming that $p_{\theta}(X_{t_{k-1}}|X_{t_{k}})$ factorizes as $\prod_{i=1}^{L}p_{\theta}(X_{t_{k-1}}^{i}|X_{t_{k}})$, following previous works (Nie et al., 2025; Sahoo et al., 2024; Lou et al., 2023).
Training process. At each training step, given a clean sequence $X_{0}$, we sample a random timestep $t\in[0,1]$ and obtain $X_{t}\sim q(X_{t}|X_{0})$ through the forward process defined in Equation 3. The loss is then computed using Equation 1 from Section 2.1.
In this section, we have documented the standard training and inference process for typical MDMs. Our modality-aware masking design introduces several modifications to the above processes, which are described in Section 3.1.2 of the main paper. Additional details are provided in Appendix A.3.
A.2 Elastic-MoT Architecture
In this section, we document the detailed design of the Elastic-MoT architecture described in Section 3.1.1. As discussed in the main paper, the proposed Elastic-MoT architecture has two key differences compared to standard MoT: a generation branch with variable size and decoupled joint attention in the later layers.
Variable-sized generation branch. In standard MoT models such as BAGEL (Deng et al., 2025), the generation branch is initialized as an exact copy of the understanding branch. For models in the 7–10B scale, this leads to a substantial increase in parameter count and compute overhead, limiting the scalability of MoT models. Motivated by the success of many medium-sized, high-quality text-to-image generation models, we explore using a smaller generation branch in the Elastic-MoT design. Since we still want modalities to interact with each other through the joint attention mechanism, it is important to keep the dimensions of the query and key vectors consistent. We provide a detailed breakdown of parameter sizes in Table 6. To initialize the generation branch with smaller dimensions than the understanding branch, we truncate the weights of the understanding branch and copy them to the generation branch.
Table 6: Comparison of understanding (Und) branch and generation (Gen) branch configurations. The projection sizes are in the format [output_size, input_size].
Und Branch Gen Branch
Attention
norm 4096 2048
q_proj_size [4096, 4096] [4096, 2048]
k_proj_size [4096, 4096] [4096, 2048]
v_proj_size [4096, 4096] [4096, 2048]
attn_out [4096, 4096] [4096, 2048]
MLP
norm 4096 2048
input_size 4096 2048
hidden_size 12288 8192
output_size 4096 2048
Figure 6: Activated parameters of Lavida-O under different task settings. Elastic-MoT Design allow Lavida-O to dynamically loads its parameters depending on the tasks. For understanding-only tasks, we only load the 8B generation branch. For text-to-image generation tasks, we load the first $M=16$ layers of the understanding branch, which consists of 4B parameters, and the full 2.4B generation branch. For interleaved tasks, we load all 2.4B+8B parameters.
Decoupled attention. In standard MoT, understanding and generation tokens can interact with each other in all $N$ transformer layers through the joint attention mechanism. We decouple the attention in the last $K$ layers and only allow tokens of the same type to interact with each other. In the first M=NKM=N-K layers, all tokens can still interact with each other as in the standard MoT architecture. This design is motivated by two factors. First, it prevents text and image tokens from interfering with each other’s representations in the later stages of generation. Second, and more importantly, it allows us to load only 4B out of 8B parameters for text-to-image generation tasks, greatly improving the scalability of pretraining while also reducing compute cost at inference time. We visualize the activated parameters for different tasks in Figure 6. For understanding-only tasks, we activate only the understanding branch in all $N=M+K$ layers. For generation-only tasks, we activate the understanding branch in the first $M$ layers and the generation branch in all $N=M+K$ layers. For interleaved tasks with both text and image outputs, we activate all parameters. In our setup, we choose $M=K=16$, yielding $N=32$ layers in total.
A.3 Modality-Aware Masking
In this section, we provide details of the changes to the training and sampling process introduced by modality-aware masking, as described in Section 3.1.2. Recall that in adapting the MoT architecture for MDMs, one of the main challenges is routing tokens. In particular, while we can easily decide which branch should process unmasked tokens based on whether they are image VQ tokens or text tokens, it is difficult to make such decisions for masked tokens, especially in interleaved generation tasks where the final output contains both images and text. Modality-aware masking addresses this problem by processing all tokens with the understanding branch by default and dynamically deciding when and where to invoke the generation branch during the sampling process.
Sampling Process. For convenience, we denote masked tokens that will be processed by the understanding branch as $M_{und}$ and masked tokens that will be processed by the generation branch as $M_{gen}$. With this distinction, the routing policy becomes simple: all text tokens plus $M_{und}$ are processed by the understanding branch, while all image VQ tokens plus $M_{gen}$ are processed by the generation branch. We introduce a special text token [exp] to indicate when an image should be generated. When a [exp] token is generated in the unmasking process, it is automatically replaced with a sequence of $M_{gen}$ tokens. The number of $M_{gen}$ tokens representing each image is determined by prespecified output resolution. These tokens are then processed by the generation branch in subsequent rounds. For example, each $1024\times 1024$ image is represented by 1024 VQ tokens. This process is documented in Algorithm 1 and illustrated in Figure 7 (Left).
Algorithm 1 Interleaved Generation with Modality-Aware Masking
1:Initial Generation Length $L$, discrete timestamps $0=t_{0}<t_{1}<\dots<t_{K}=1$, prompt $C$
2:Initialize $t\leftarrow K$
3:Initialize $X_{t}^{1:L}\leftarrow M_{und}$
4:for $i=T$ to 1 do
[…]
6:  if a [exp] token is generated in $X_{t_{i-1}}$ then
7:   Replace it with a sequence of $M_{gen}$ tokens
8:    // These $M_{gen}$ will be routed to the generation branch in subsequent rounds
9:  end if
10:end for
11:return Fully unmasked sequence $X_{0}$
Figure 7: Training and inference with modality-aware masking. We visualize the sampling process with modality-aware masking on the left and the training process on the right. During the training, the loss is applied on either $X_{0}$ or $X_{0}^{\prime}$ depending on the value of $t$ with respect to $t_{\text{exp}}.$
Training. A consequence of modality-aware masking is that the partially masked sequence $X_{t}$ will have varying length depending on $t$, making the loss described in Equation 1 not directly applicable. In particular, when sampling from the forward process $q(X_{t}\mid X_{0})$, there is a special timestep $t_{\text{exp}}$ at which a sequence of VQ image tokens is collapsed into a single [exp] text token. As illustrated in Figure 7 (Right), when $t<t_{\text{exp}}$, $X_{t}$ has a shorter sequence length than $X_{0}$. To apply the loss properly, we construct a new sequence $X_{0}^{\prime}$ by collapsing all sequences of image VQ tokens into [exp] tokens in $X_{0}$. We then modify the loss in Equation 1 to the following:
$\mathcal{L}_{\text{MDM}}=-\mathbb{E}_{t,X_{0},X_{t}}\!\left[\frac{1}{t}\sum_{\{i\mid X_{t}^{i}=[M]\}}\log p_{\theta}(\hat{X}_{0}^{i}\mid X_{t})\right],$ (6)
$\text{where }\hat{X}_{0}=\begin{cases}X_{0},&\text{if }t\in(t_{\text{exp}},1)\\ X_{0}^{\prime},&\text{if }t\in(0,t_{\text{exp}})\end{cases}$ (7)
This change is also highlighted in blue in Figure 7 (Right).
Understanding-Only and Generation-Only Tasks. We activate modality-aware masking only for interleaved tasks, since these require both the understanding and generation branches in our Elastic-MoT architecture. For computational efficiency, we do not use modality-aware masking for understanding-only tasks such as image captioning, or for generation-only tasks such as text-to-image generation (without planning and reflection). This allows us to best utilize the flexibility of Elastic-MoT and avoid loading unnecessary model parameters.
Figure 8: Effect of Universal Text Conditioning. On the left side, we visualize the text format used in Universal Text Conditioning. On the right side, we visualize generation results under different choices of universal text conditioning.
A.4 Universal Text Conditioning
Universal text conditioning is inspired by the micro-conditioning approach (Podell et al., 2023) employed in many text-to-image models. These models interoperate special conditioning embeddings to incorporate non-text conditions such as the original image resolution or aesthetic score. Since Lavida-O is a unified model with mathematical reasoning capabilities, we can represent these conditions directly as plain text. In particular, we include source image resolution, crop coordinates, aesthetic scores (Schuhmann, 2022), and HPS scores (Wu et al., 2023), following existing works (Podell et al., 2023; Bai et al., 2024). Additionally, we incorporate luminance (brightness) and contrast to give users greater control over the generated images. Each condition is represented as a simple string of the form “[KEY] : [VALUE]”. During training, each condition is randomly dropped with some probability. At inference, users may specify all conditions or only a subset.
This design is illustrated in Figure 8. By modifying these universal text conditioning parameters at inference time, users can flexibly control various image properties such as brightness. Notably, when brightness and contrast are set to very high values, the generated images become highly stylized in order to satisfy the constraints.
A.5 Stratified Random Sampling
Figure 9: Visualization of different sampling processes. We compare the unmasking order of the stratified sampler, Halton sampler, and uniform random sampler. Uniform random sampler produces the least desirable spatial pattern, with many unmasked tokens clustered together. Halton sampler is less ideal than stratified sampler because it does not guarantee perfectly stratified coverage. For example, when the number of unmasked tokens is 4, the upper-right quadrant remains unoccupied.
In this section, we provide detailed descriptions of the stratified random sampling process introduced in Section 3.2. In the vanilla sampling process described in Equation 5, each token is unmasked independently. In practice, this often leads to suboptimal generation quality. Instead of unmasking tokens randomly, several works adopt alternative sampling strategies in which the unmasking order of tokens is determined by heuristics such as the model’s confidence at each token position (Nie et al., 2025; Ye et al., 2025a; Chang et al., 2022).
In image generation, tokens with high confidence are frequently adjacent to one another. As a result, confidence-based unmasking tends to reveal many adjacent tokens in a single step. Since tokens that are spatially adjacent are often highly correlated, this violates the independence assumption $p_{\theta}(X_{t_{k-1}}|X_{t_{k}})=\prod_{i=1}^{L}p_{\theta}(X_{t_{k-1}}^{i}|X_{t_{k}})$ stated in Section A.1. To address this, we design a stratified sampling process that ensures the unmasked tokens are spatially dispersed. Specifically, we enforce that the first 4 unmasked tokens occupy the four quadrants of the image; the first 16 unmasked tokens occupy all 16 subregions obtained by dividing the image into a $4\times 4$ grid; and so forth. The algorithm is formally described below:
Algorithm 2 Stratified Unmasking Order
1:Image size $N\times N$
2:a list $\mathcal{O}$ of coordinates $(i,j)$ indicating unmasking order
3:Initialize an empty list $\mathcal{O}$
4:for $d=1,2,\dots,\log_{2}N$ do
5:  Partition the image into $2^{d}\times 2^{d}$ grid cells
6:  for each grid cell $g$ in random order do
7:   if $\mathcal{O}\cap g=\text{\O }$ then
8:     Sample $(i_{g},j_{g})$ uniformly within cell $g$
9:     Append $(i_{g},j_{g})$ to $\mathcal{O}$
10:   end if
11:  end for
12:end for
13:return $\mathcal{O}$
Our design is inspired by the stratified sampling process commonly used in numerical integration and computer graphics. It also follows a similar motivation to the recent Halton mask scheduler, which uses the low-discrepancy Halton sequence to ensure that unmasked tokens are spatially dispersed (Besnier et al., 2025). We illustrate the differences among stratified sampling, Halton sampling, and uniform random sampling in Figure 9. As shown in the figure, uniform random sampling produces the least desirable spatial pattern, with many unmasked tokens clustered together. Compared with our proposed stratified sampling process, Halton sampling is less ideal because it does not guarantee perfectly stratified coverage. For example, when the number of unmasked tokens is 4, the upper-right quadrant remains unoccupied. The benefits of stratified sampling are also reflected in FID scores, which we document in Section B.3.
A.6 Object Grounding with Coordinate Quantization
In this section, we provide detailed descriptions of Lavida-O’s design for object grounding tasks. Given an image and a referring expression describing an object, the grounding task requires locating the described object in the image by predicting its bounding box coordinates. In autoregressive vision-language models such as Qwen2.5-VL (Bai et al., 2025), bounding boxes are represented as plain text strings, such as “[123, 232, 300, 1021]”. At inference, the coordinates are generated sequentially from left to right. This design has several limitations. First, since the model only sees a padded and resized image, it is difficult for the model to predict absolute pixel coordinates that depend on the original resolution of the input image. Second, the sequential generation order is slow and inefficient.
To address these issues, we normalize the bounding box coordinates and quantize them into discrete bins. Specifically, given an image of size $H\times W$, we first pad it to a square image of size $D\times D$, where $D=\max(H,W)$, and normalize the bounding boxes in the padded image to the range [0,1] by dividing the raw pixel coordinates by $D$. This step makes the coordinates independent of the original input resolution. We then round each coordinate into 1025 bins representing $\frac{0}{1024},\frac{1}{1024},\frac{2}{1024},...,\frac{1024}{1024}$ and represent them with special tokens. This reduces the number of tokens needed to represent each bounding box to exactly 4. Finally, since Lavida-O is a masked diffusion model with a bi-directional attention mask and parallel decoding capabilities, we can predict multiple bounding boxes simultaneously. For example, if we want to obtain the bounding boxes of both “a cute dog” and “a boy,” we can initialize a text sequence “a cute dog [m][m][m][m]; a boy [m][m][m][m]” and perform parallel unmasking of multiple bounding box coordinates. This design is illustrated in Figure 10.
Figure 10: Coordinate Quantization. We normalize bounding box coordinates into the range [0,1] and discretize them into 1025 bins. This ensures that each bounding box is represented by exactly 4 tokens, allowing efficient parallel decoding of multiple bounding boxes in a single step.
A.7 Reflection and Planning
The unique advantage of unified understanding and generation models is that they can leverage their understanding capabilities to improve generation results. Several works on unified models show that simple joint training on a combination of understanding and generation tasks improves performance on generation tasks (Xie et al., 2024; Deng et al., 2025), particularly in instruction-following capabilities. Lavida-O pushes this paradigm further by introducing two explicit mechanisms to exploit understanding capabilities: planning and reflection.
Figure 11: Interleaved Generation with Planning and Reflection. We provide visual examples of interleaved generation, including text-to-image generation with planning (Top), text-to-image generation with reflection (Middle), and image editing with planning (Bottom). We always enable planning during the reflection process. The layout traces is omitted in the middle figure for clarity and better presentation.
Planning. To improve prompt-following capabilities in text-to-image generation, we ask the model to first generate a layout design of objects, which consists of (object, bounding box) pairs, before generating the final image. Such interleaved generation is achieved through the modality-aware masking process described in Section A.3. We illustrate this process in Figure 11 (Top). As shown, planning enables Lavida-O to follow challenging and unintuitive prompts, such as “a horse above an astronaut.”
Similarly, we can adopt planning for image editing tasks. Given an input image and an edit instruction, the model can first leverage its grounding capabilities to identify the regions that need to be edited before generating the edited image. This process is illustrated in Figure 11 (Bottom).
Reflection. We can improve text-to-image generation performance by leveraging Lavida-O’s understanding capability to achieve self-critique and iterative self-improvement. Given an input prompt, the model first generates an image, then performs a self-critique step to evaluate whether the generated image matches the prompt. If it does, the generation process terminates. Otherwise, the model generates a revised image and attempts to fix the identified issues. This cycle is repeated until an image passes the self-critique process or the maximum number of rounds is reached. At each round, we also invoke the planning capability. Since Lavida-O’s context length is limited to 8192 tokens, we truncate the history when necessary to include at most three rounds. This process is illustrated in Figure 11 (Middle).
Similar designs have been explored for generation-only models in the context of inference-time scaling, such as Reflect-DiT (Li et al., 2025b) and ReflectionFlow (Zhuo et al., 2025). However, unlike these works, which require an external vision-language model as a reward model, Lavida-O uniquely unifies layout planning, self-critique, and iterative self-improvement in a single model through a unified generation process.
Appendix B Additional Experiment Details and Results
In this section, we document experiment details for better reproducibility, including data pipeline, training hyperparameters, and compute cost. In addition, we also provide additional experimental results on the effectiveness of various design choices employed by Lavida-O, such as the Elastic-MoT design, stratified sampling, and the data pipeline.
B.1 Setup
Pretrained Weights. We use LaViDa (Li et al., 2025a) to initialize the understanding branch and semantic encoder. For the VQ-encoder, we adopt the encoder of Meissonic (Bai et al., 2024). The image generation branch is initialized from the truncated weights of the understanding branch, as described in Section A.2.
Data Pipeline. Unlike many frontier models, our model does not make use of any proprietary images or documents. Our training data consists of the following components:
  • A: Text-to-Image Pairs. We source data from LAION-2B (Schuhmann et al., 2022) and COYO-700M (Byeon et al., 2022). We additionally include SA-1B (Kirillov et al., 2023), JourneyDB (Sun et al., 2023), BLIP3o-60k (Chen et al., 2025a), and ShareGPT4o-Image (Chen et al., 2025b). Each dataset is heavily filtered to remove NSFW prompts, low CLIP scores (Radford et al., 2021), low aesthetic scores (Schuhmann, 2022), and low-resolution images. This results in 200M images in our final mix. Where available, we use captions generated by VLMs instead of raw alt-texts. These captions are sourced from existing works including Recap-LAION, Recap-COYO (Wu et al., 2024), and BLIP-3o (Chen et al., 2025a).
  • B: Image-level Understanding Data. We include LLaVA-OneVision (Li et al., 2024a), Open-LLaVA-Next (Chen & Xing, 2024), MAmmoth-VL (Guo et al., 2024), and VisualWebInstruct (Jia et al., 2025).
  • C: Region-level Understanding Data. We include GranD (Rasheed et al., 2024) and RefCOCO (Kazemzadeh et al., 2014).
  • D: Image Editing Data. We include ShareGPT4o-Image (Chen et al., 2025b), GPT-Edit-1.5M (Wang et al., 2025b), and the image editing subset of UniWorld-V1 (Hu et al., 2022).
  • E: Interleaved Planning and Reflection Data. For planning data, we manually construct a layout dataset by running an open-vocabulary object detector, GroundingDino-L (Liu et al., 2024b), on the outputs of image generation and editing datasets, including BLIP-3o (Chen et al., 2025a), ShareGPT4o-Image (Chen et al., 2025b), and GPT-Edit-1.5M (Wang et al., 2025b). For reflection data, we leverage existing datasets including ReflectDiT (Li et al., 2025b) and ReflectionFlow (Zhuo et al., 2025).
Training Setup. Training consists of three stages. In the first stage, we extend LaViDa to region-level tasks such as grounding. In the second stage, we perform large-scale pretraining on text-to-image generation tasks. In the final stage, we jointly train the model on a mix of understanding, generation, and interleaved tasks. We document the training hyperparameters, datasets used, active parameter count, and other relevant details in Table 7.
In addition, we implement a dataset mix scheduler that dynamically adjusts the sampling weight of each dataset throughout training to address data imbalance. Specifically, we assign a high weight to new capabilities at the start of each training stage and gradually decay the weight over time. For example, in Stage 1 we have fewer than 1M grounding samples but over 10M image-level understanding samples. To enable efficient acquisition of grounding capability while preventing overfitting, we set the grounding-to-understanding ratio to 3:1 initially, which is gradually decreased to 1:3. We provide further analysis of the scheduler in Section B.4.
Table 7: Training configurations across three stages. We use letters A-E to represent different dataset following Section B.1.
Stage 1 Stage 2 Stage 3
Learning Rate $5\times 10^{-6}$ $1\times 10^{-4}$ $2\times 10^{-5}$
Steps 80k 400k 100k
$\beta_{1}$ 0.99 0.99 0.99
$\beta_{2}$ 0.999 0.999 0.999
optimizer AdamW AdamW AdamW
Dataset Used B,C A A,B,C,D,E
Loaded Parameters 8B 6.4B 10.4B
Trainable Parameters 8B 2.4B 10.4B
Und. resolution 384 $\times\{(1,3),(2,2)\}$ 384 $\times\{(1,3),(2,2)\}$ 384 $\times\{(1,3),(2,2)\}$
Gen. resolution - 256 $\rightarrow$ 512 $\rightarrow$ 1024 1024
Semantic Encoder Trainable Not Loaded Trainable
VQ Encoder Not Loaded Loaded Loaded
Gen. Branch Not Loaded Trainable Trainable
Und. Branch Trainable Partially Loaded Trainable
B.2 Ablation Studies on Elastic-MoT Design
In this section, we report ablation results of the Elastic-MoT design, including the size of the generation branch and the number of joint attention layers.
Size of Generation Branch. We report the performance of Lavida-O with different sizes of the generation branch during text-to-image pretraining (Stage 2) in Table 8. Results are obtained after 50k training steps with a global batch size of 1024. We also document the maximum per-GPU batch size, gradient accumulation steps, and training latency to measure efficiency. Results show that models of different sizes achieve comparable performance after 50k steps. Smaller models (1B, 2B) converge slightly faster and achieve marginally higher performance than larger models (4B, 8B). In terms of latency, smaller models are considerably faster. The 2B model achieves the best balance between performance and efficiency, attaining the highest GenEval and DPG scores while being $3.17\times$ faster.
Table 8: Comparison of different model sizes on GenEval, DPG, and training efficiency. We report the performance of Lavida-O with different sizes of the generation branch during the text-to-image pretraining (Stage-2) after 50k training steps. We also report the per-GPU batch size and training latency.
Architecture Performance Efficiency
[…]
4B+1B 1536 0.56 60.8 16 1 1.98
4B+2B 2048 0.57 63.1 16 1 3.67
4B+4B 3072 0.48 55.3 8 2 8.42
4B+8B 4096 0.55 58.6 8 2 11.64
Number of Joint Attention Layers. To study the effect of varying the number of joint attention layers, we conducted two ablation experiments. The first experiment was performed during Stage 2 pretraining. We started with the Stage 1 checkpoint with $N=32$ layers in the understanding branch and fixed the generation branch size to 4B. We then varied $M$, the number of layers with joint attention, among {8,16,24,32}. The number of non-joint layers, $K$, is automatically determined by K=NMK=N-M. Results after 100k training steps are shown in Table 9. Among the four choices, $M=\{16,24,32\}$ yield comparable performance, while $M=8$ shows a substantial drop. This suggests that a sufficient number of joint attention layers is necessary for strong text-to-image performance, but additional layers beyond a threshold provide little benefit. Training latency also decreases when $M$ is smaller (i.e., larger $K$), since fewer joint layers need to be loaded. $M=16$ achieves the best balance of speed and performance.
We conducted a second experiment in Stage 3, where interleaved generation and editing tasks may benefit more from joint attention. Starting from a Stage 2 checkpoint pretrained with $M=16$ layers for 400k steps, we trained for 50k steps under $M=\{16,24,32\}$. Results show two key observations: (1) text-to-image tasks converge faster than image-editing tasks, reaching near-final performance after 50k steps, while editing tasks lag behind; (2) increasing $M$ does not significantly improve performance, even for interleaved editing. This may be due to token interference at later layers or the Stage 2 model being optimized with only 16 joint layers. Due to compute constraints, we were unable to retrain Stage 2 with alternative values of $M$. Nevertheless, keeping $M=16$ is a reasonable choice given our setup. Finally, in Stage 3 the efficiency difference is less pronounced, since all 10.4B parameters must be loaded for interleaved training and inference.
Table 9: Effect of varying choices of M and K in partially-decoupled attention design. Efficiency is measured in Stage-2 training. For stage-3 training, we need to load all layers since the data contain a mix of text, image, and interleaved generation tasks.
M K Pretraining (Stage-2) SFT (Stage-3) Efficiency
[…]
8 24 0.57 69.3 - - - 2.45
16 16 0.63 75.0 0.89 83.2 3.66 3.67
24 8 0.63 73.3 0.81 83.0 3.60 4.12
32 0 0.61 71.2 0.85 83.2 3.55 5.20
Weight Initialization. We initialized the 2.4B generation branch with truncated weights from the understanding branch (Section A.2). We also explored initializing from scratch. Figure 12 shows validation loss during the first 20k steps of Stage 2. Truncated initialization converges faster and yields lower loss.
Figure 12: Effect of truncated initialization. Validation loss comparison of truncated initialization vs. training from scratch during Stage 2. Truncated initialization converges faster and achieves lower loss.
B.3 Ablation Studies on Stratified Sampling
We compared image generation quality under different sampling strategies on the MJHQ-30k dataset (Li et al., 2024b) with 64 sampling steps. We evaluate the proposed stratified sampler against confidence-based sampling (Chang et al., 2022), uniform random sampling, and Halton sampling (Besnier et al., 2025). Results are reported in Table 10. The stratified sampler achieves the best performance.
Table 10: Performance of Different Samplers in Text-to-Image Generation Tasks. We report the FID scores on MJHQ-30K dataset using different samplers. The proposed stratified sampler achieves the best outcome.
Method FID-30k $\downarrow$
Confidence 11.42
Uniform 8.22
Halton 7.38
Stratified 6.68
B.4 Ablation Studies on Data Pipeline
Effect of Task Scheduler. To study the effect of the dataset scheduler described in Section B.1, we compare three dataset mixing strategies in Stage 1 training. The goal of Stage 1 is to equip LaViDa with region-level understanding capabilities such as grounding. At this stage, training data includes fewer than 1M grounding samples but over 10M image-level understanding samples. To mitigate imbalance, we employ a scheduler that dynamically adjusts the sampling weights for new (grounding) and existing (image-level) capabilities. Each batch is drawn from a single dataset. For example, when New:Old=1:3, on average $\tfrac{1}{4}$ of batches contain grounding data and $\tfrac{3}{4}$ contain image-level data.
We initialize the ratio as New:Old=3:1 and gradually reduce it to 1:3. We compare against fixed ratios of 1:3 and 3:1, reporting results after 20k steps in Table 11. Fixing New:Old=1:3 under-trains grounding, while fixing New:Old=3:1 improves grounding but causes forgetting on image-level understanding. By contrast, the dynamic scheduler achieves strong performance on both. Notably, it even outperforms the fixed 3:1 setup on image-level understanding, suggesting it also mitigates overfitting caused by the small grounding dataset.
Table 11: Comparison of different task scheduling during Stage 1 Training. We compare the performance under different dataset sampling weights of new capabilities (grounding) and old capabilities (image-level understanding). We explored two fixed sampling ratio 1:3 and 3:1 for New:Old. For the dynamic scheduler, the New:Old ratio is initialized as 3:1 and gradually decreased to 1:3.
Method New Capabilities Existing Capabilities
RefCOCO RefCOCO+ RefCOCOg MME ChartQA ScienceQA
New:Old = 1:3 83.2 74.6 78.3 449 72.6 84.3
New:Old = 3:1 88.8 82.4 85.7 349 65.0 75.8
Dynamic 92.0 86.9 89.3 436 73.4 86.4
Does understanding data help generation tasks? To examine whether incorporating understanding data benefits generation, we experimented with removing all grounding data from Stage 3. Results are shown in Table 12. Even without invoking explicit planning, grounding data improves text-to-image and editing tasks, suggesting synergy between the two. When planning is enabled, the gains are further amplified.
Table 12: Effect of Grounding Data in Stage 3 Training. To study the synergy between understanding and generation tasks, we explored removing object grounding in Stage 3 Training. This leads to worse overall performance, suggesting jointly training on both tasks imporve the generation performance.
Method GenEval DPG ImageEdit
w/o grounding data 0.74 82.0 3.60
w/ grounding data 0.77 81.8 3.71
+ planning 0.85 82.9 3.80
B.5 Ablation Studies on Reflection and Planning
Breakdown of Performance Improvements. We provide a detailed breakdown of planning and reflection gains. Table 13 shows results on GenEval. Planning yields large improvements in object positioning (+0.19), while reflection additionally improves counting and attribution.
On Image-Edit (Table 14), planning improves adding/removing objects, subject actions, and hybrid instructions. The largest gains are in action (+0.50) and hybrid (+0.16). However, global edits (e.g., style, background) degrade slightly, as these tasks are less aligned with grounding. A promising direction for future works is to let the model dynamically decide whether to invoke planning.
Table 13: Breakdown of performance improvements on GenEval Dataset. We report the improvements of the planning and reflection mechanism on each category of the text-to-image generation tasks from GenEval Dataset.
Single Two Position Counting Color Attribution Overall
Baseline 0.99 0.85 0.65 0.71 0.86 0.58 0.77
+Planning 0.99 0.94 0.84 0.75 0.90 0.68 0.85
$\Delta$ vs. Baseline = +0.09 +0.19 +0.04 +0.04 +0.10 +0.08
+Reflection 1.00 0.95 0.89 0.85 0.90 0.74 0.89
$\Delta$ vs. Baseline +0.01 +0.10 +0.24 +0.14 +0.04 +0.16 +0.12
Table 14: Breakdown of performance improvements on Image-Edit Dataset. We report the improvements of the planning mechanism on each category of the image editing tasks from Image-Edit Dataset.
Model Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
Baseline 4.04 3.62 2.01 4.39 3.98 4.06 4.82 2.94 3.54 3.71
+ Planning 4.11 3.67 2.04 4.40 4.05 4.00 4.75 3.10 4.04 3.80
$\Delta$ vs. Baseline +0.07 +0.05 +0.03 +0.01 +0.07 -0.06 -0.07 +0.16 +0.50 +0.09
Effect of Inference-time Scaling. We evaluate reflection scaling by varying $N$, the maximum number of images generated per prompt. Table 15 shows results. Even one reflection step ($N=2$) improves performance. Gains saturate at $N=8$, with little benefit beyond. Latency grows sublinearly with $N$ since simple prompts often trigger early stopping. For example, when $N=20$, the model may obtain a satisfying output and terminate the generation process after generating just two images.
Table 15: Performance and Latency at different numbers of reflection rounds NN. When $N=1$, we only perform planning.
Num. of Reflection Rrounds N=1 N=2 N=4 N=8 N=12 N=16 N=20
GenEval Score $\uparrow$ 0.848 0.864 0.875 0.882 0.890 0.886 0.886
Latency (s/image) $\downarrow$ 27.2 32.6 39.3 47.1 53.4 58.3 62.2
B.6 Speed–Quality Tradeoff
A key advantage of masked diffusion models over autoregressive models is the speed–quality tradeoff enabled by parallel decoding. We study this in the unified setting by evaluating Lavida-O on MJHQ-30k text-to-image generation (Li et al., 2024b), RefCOCO grounding (Kazemzadeh et al., 2014), and MathVista reasoning (Lu et al., 2023).
For MJHQ and RefCOCO, we vary the number of diffusion steps. For MathVista, we employ Fast-DLLM (Wu et al., 2025a), which adaptively unmasks multiple tokens per step. The tradeoff is controlled via its threshold hyperparameter. Results are shown in Figure 13. For MJHQ we report FID (lower is better), for RefCOCO Precision@0.5 (higher is better), and for MathVista accuracy (higher is better).
We compare against several baselines: Flux (Labs, 2024) on T2I, Qwen2.5-VL-7B (Bai et al., 2025) on grounding, and Qwen2.5-VL/Open-LLaVA-Next-8B (Chen & Xing, 2024) on reasoning. Lavida-O achieves faster inference and stronger quality on image generation and grounding. For grounding, it reaches up to $6.8\times$ speedup while surpassing Qwen2.5-VL-7B in precision. On MathVista, while less accurate than state-of-the-art AR models, Lavida-O is much faster, and still stronger than popular AR baselines such as Open-LLaVA-Next-8B. Performance also exceeds the base LaViDa (56.9 vs. 44.8).
Figure 13: Speed–quality tradeoff on generation, grounding, and reasoning. Latency (s/sample) and benchmark scores are shown. For MJHQ: FID (lower is better). For RefCOCO: Precision@0.5 (higher is better). For MathVista: accuracy (higher is better). On MathVista, the maximum generation length is capped at 256 tokens.
B.7 Additional Qualitative Results
Finally, we provide additional qualitative examples demonstrating Lavida-O’s capabilities on diverse prompts and editing instructions. Figure 14 shows text-to-image generation, and Figure 15 shows image editing results.
Appendix C Compute Cost
All experiments are conducted on 8 nodes, each equipped with 8 A100 GPUs. The total training amounts to 34.2 days measured by wall clock time, or 53k GPU hours.
Appendix D Limitations
In this section, we discuss several limitations of Lavida-O.
Text Rendering. Since the image generation branch is trained from scratch and we did not explicitly include datasets for text rendering, Lavida-O ’s capability to render and edit text is very limited. We also find that the VQ image tokenizer we use can not faithfully reconstruct small texts. We aim to address this issue in future works by incroprating additional text rendering data and finetune the VQ image tokenizer on screenshots of documents.
Pixel Shift. Our image editing datasets, such as GPT-Image-Edit-1.5M (Wang et al., 2025b) contains images distilled from generative models like GPT-4o, which is know to have “pixel shift” problems. Specifically, even if the instruction only requires editing a specific region, the rest of image may still experience small but noticeable changes. As a consequence, Lavida-O inherit this problem. We aim to mitigate this by obtaining more clean and high-quality image-editing data.
Math Reasoning. The focus of Lavida-O is to build a unified multi-modal MDMs capable of both understanding and generation tasks. While its math reasoning capabilities has improved from the base model LaViDa thanks to additional training, there remain a considerable gap when compared against state-of-the-art models. We left further improvements on math reasoning tasks to future works.
Hallucination. Like all generative models, our model may suffer from hallucination problems at inference. We recommend treating model output with caution instead of blind trust.
Appendix E Boarder Impact
Lavida-O has strong text-to-image generation capabilities and image-editing capabilities, which may be abused to create various harmful and offensive content. We strongly caution the community against such usecases. Additonally, our model may inherit the biases embedded in the base model LaViDa, as well as biases incorporated in the images and texts of the training data. Our model is intended to be used by researchers to build a strong diffusion model for multi-modal applications and explore methods of building future multi-modal foundational models. We do not recommend it be used for any other purposes.
Figure 14: Qualitative examples of text-to-image generation. We provide additional examples of text-to-image generation outputs on diverse prompts.
Figure 15: Qualitative examples of image editing. We provide additional examples of image editing outputs on diverse instructions.